In a book of 520 pages, 390 typo

-graphical errors occur. Assuming Poisson law for the

number of errors per page, find the probability that a random sample of 5 pages will

contain no error.


1
Expert's answer
2021-05-07T10:09:09-0400

The average number of typographical errors per page in the book is given by

"λ = \frac{390}{520} = 0.75"

Hence using Poisson probability law, the probability of x errors per page is given by:

"P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

= \frac{e^{-0.75} \times (0.75)^x}{x!}"

The required probability that a random sample of 5 pages will contain no error is given by:

"[P(X=0)]^5 = (e^{-0.75})^5 = e^{-3.75}"